# A Hausdorff-measure boundary element method for acoustic scattering by fractal screens

### v1

@@ 4.1 The BIE on $d$-sets in trace spaces
$\langle\mathbb{S}\Psi,\widetilde{\Psi}\rangle_{\mathbb{H}^{t_{d}}(\Gamma)\times\mathbb{H}^{-t_{d}}(\Gamma)}=-\langle\mathrm{tr}_{\Gamma}g,\widetilde{\Psi}\rangle_{\mathbb{H}^{t_{d}}(\Gamma)\times\mathbb{H}^{-t_{d}}(\Gamma)},\qquad\widetilde{\Psi}\in\mathbb{H}^{-t_{d}}(\Gamma),$ (49)
@@
The definition and mapping properties of $\mathrm{tr}_{\Gamma}$ and $\mathrm{tr}_{\Gamma}^{*}$, noted in §2.4, combined with the representation (50), enable us to extend the domain of $\mathbb{S}$ to $\mathbb{H}^{-t}(\Gamma)$, for $t_{d}<t<2t_{d}$, or restrict it to $\mathbb{H}^{-t}(\Gamma)$, for $0<t<t_{d}$, as stated in the following key lemma.
@@
Let $\Omega\subset\Gamma_{\infty}$ be any bounded open set containing $\Gamma$, so that (see §2.4) $H^{s}_{\Gamma}$ is a closed subspace of $\widetilde{H}^{s}(\Omega)$ for every $s\in\mathbb{R}$. The claimed mapping property of $\mathbb{S}$ follows from (50) since $\mathrm{tr}_{\Gamma}:H^{s}(\mathbb{R}^{n})\to\mathbb{H}^{t}(\Gamma)$ and $\mathrm{tr}_{\Gamma}^{*}:\mathbb{H}^{-t}(\Gamma)\to H^{-s}_{\Gamma}\subset\widetilde{H}^{-s}(\Omega)$ are continuous for s>(nd)/2s>(n-d)/2 (i.e. $t>0$), with $s$ and $t$ related by (10), and since the mapping $\phi\mapsto\gamma^{\pm}((\sigma\mathcal{S}\phi)|_{U^{\pm}}):\widetilde{H}^{s}(\Omega)\to H^{s+1}(\Gamma_{\infty})$ is continuous for $s\in\mathbb{R}$ (e.g., [11, Thm 1.6]). ∎
@@
We recall the following result from [33, Prop. 4.7], which quotes [38]. Suppose $E_{j}$ and $F_{j}$ are Banach spaces, $E_{1}\subset E_{0}$ and $F_{1}\subset F_{0}$ with continuous embeddings, and $T:E_{j}\to F_{j}$ is a bounded linear operator for $j=0,1$. Let $E_{\theta}=(E_{0},E_{1})_{\theta}$ be defined by complex interpolation, and similarly $F_{\theta}$, so $T:E_{\theta}\to F_{\theta}$ and is bounded, for $0<\theta<1$. Assume that $T:E_{\theta_{0}}\to F_{\theta_{0}}$ is invertible for some $\theta_{0}\in(0,1)$. Then $T:E_{\theta}\to F_{\theta}$ is invertible for $\theta$ in a neighbourhood of $\theta_{0}$.
@@ 5 The Hausdorff BEM
$a(\phi_{N},\psi_{N})=-\langle g,\psi_{N}\rangle_{H^{1/2}(\Gamma_{\infty})\times H^{-1/2}(\Gamma_{\infty})},\qquad\psi_{N}\in V_{N}.$ (56)
@@
The well-posedness of (56) follows from the Lax–Milgram lemma and the continuity and coercivity of $)S:H^{-1/2}_{\Gamma}\to(\widetilde{H}{}^{1/2}(\Gamma^{c}$$)^{\perp}$. Furthermore, by Céa’s lemma (e.g., [40, Theorem 8.1]) we have the following quasi-optimality estimate, where $C_{a},\alpha>0$ are as in (41):
@@ 5.3 Inverse estimates and conditioning
The first claim of the following corollary follows immediately from the above result. The second (cf. [21, Thm. 1.4], [18], [17, Thm. 3.2.4]) combines our earlier results with standard Aubin-Nitsche lemma arguments.
@@ 5.4 Numerical quadrature and fully discrete error estimates
For $i\neq j$ the integrand in (93) is continuous, because, by assumption, the components of $\Gamma$ are disjoint. In that case, to evaluate (93) we use the composite barycentre rule of [25, Defn 3.5]. In our context, this means partitioning the BEM elements $\Gamma_{{\boldsymbol{m}}(i)}$ and $\Gamma_{{\boldsymbol{m}}(j)}$, which have diameter approximately equal to $h$, into a union of self-similar subsets of a possibly smaller approximate diameter $0<h_{Q}\leq h$, writing the integral in (93) as a sum of integrals over the Cartesian products of these subsets, and approximating these integrals by a one-point barycentre rule. Specifically, we approximate
@@
is the barycentre of $\Gamma_{\boldsymbol{n}}$ with respect to the measure $\mathcal{H}^{d}$. (Note that $x_{\boldsymbol{n}}$ is not necessarily an element of $\Gamma_{\boldsymbol{n}}$.) The similarities can be written as $s_{m}(x)=\rho_{m}A_{m}x+v_{m}$, for some $v_{m}\in\mathbb{R}^{n}$ and some orthogonal $A_{m}\in\mathbb{R}^{n\times n}$ ($A_{m}=\pm 1$ in the case $n=1$), and then, writing $\boldsymbol{n}=(n_{1},\ldots,n_{\ell})$, we find (see [25, Prop. 3.3]) that
@@
For $i=j$ the integral (93) is singular. To evaluate it we adopt the singularity subtraction approach of [25, §5], writing
@@
The integrals appearing in (97) all have continous integrands and can be evaluated using the composite barycentre rule (as in [25, Eqn (47)]), viz.
@@
Assuming the data $g$ is sufficiently smooth, the right-hand side entries (61) are regular single integrals and can be evaluated by a single integral version of the composite barycentre rules described for double integrals above, as in [25, Defn 3.1]). Explicitly, we use the approximation
@@
The following quadrature error estimates follow from results in [25], specifically [25, Prop. 5.2, Thms 3.6(iii), 5.7 & 5.11]. Following the terminology of [25], we say an IFS attractor $\Gamma$ is hull-disjoint if the convex hulls $\mathrm{Hull}(\Gamma_{1}),\ldots,\mathrm{Hull}(\Gamma_{M})$ are disjoint, which holds if and only if the OSC holds for some open set $O\supset\mathrm{Hull}(\Gamma)$ (cf. Lemma 2.5, the proof of which works in the same way for hull-disjointness as for disjointness). The assumption of hull-disjointness permits analysis of the singular quadrature rules in [25] by Taylor expansion. (The point here is that while the barycentre $x_{{\boldsymbol{m}}}$ of a component $\Gamma_{{\boldsymbol{m}}}$ may not lie in $\Gamma_{{\boldsymbol{m}}}$, it must lie in $\mathrm{Hull}(\Gamma_{{\boldsymbol{m}}}$).) However, numerical experiments suggest that hull-disjointness is not essential for the applicability of the quadrature rules in [25] (see [25, §6]). We also suspect that the estimates (105) and (104) may not be sharp in their $h$-dependence — see [25, Rem. 5.10] and the discussion around [25, Fig. 7]. In Remark 5.19 we describe modifications to our quadrature rule which may improve its efficiency.
@@
the first estimate in (102) follows from [25, Thm 3.6(iii)], and the second then follows from the fact that $\sum_{i=1}^{N}\mu_{{\boldsymbol{m}}(i)}=\mathcal{H}^{d}(\Gamma)$. For (ii), given $\psi_{N}\in V_{N}$ let ${\vec{\psi}}=(({\vec{\psi}})_{1},\ldots,({\vec{\psi}})_{N})^{T}\in\mathbb{C}^{N}$ denote the coefficient vector of its expansion with respect to the basis $(e^{1},\ldots,e^{N})$. By the orthonormality of the basis $(f^{1},\ldots,f^{N})$ in $\mathbb{L}_{2}(\Gamma)$, the inverse estimate (86) and the isometry property (16), given $0<t<1$ we can bound
@@
Then, to prove (103), using (65), (101), [25, Thm 3.6(iii)], (106), and the Cauchy-Schwarz inequality,
@@
For (iii), the first estimates in (104) and (105) follow from [25, Prop 5.2] for the off-diagonal terms and [25, Thms 5.7 & 5.11] for the diagonal terms. The factors of $h^{-(n+1)}$ and $h^{-n}$ come from the fact that, in the notation of [25, Thms 5.7 & 5.11], $R_{\Gamma,\mathrm{Hull}}$ equals $\operatorname{diam}(\Gamma)$ times a constant independent of $\operatorname{diam}(\Gamma)$. We note that, since we are allowing $C$ to be $k$-dependent, we can use the $k\operatorname{diam}(\Gamma)\leq c_{\rm osc}$” estimates in [25, Thms 5.7 & 5.11], since, given $k>0$, the constant $c_{\rm osc}$ can be chosen as large as is required. To derive the second estimate in (104) (a similar argument gives the second estimate in (105)), note that, with $\|\cdot\|_{\mathrm{F}}$ denoting the Frobenius norm,
$\|A-A^{Q}\|_{2}\leq\|A-A^{Q}\|_{\mathrm{F}}\leq Ch_{Q}h^{-n}\bigg(\sum_{i=1}^{N}\sum_{j=1}^{N}\mu_{{\boldsymbol{m}}(i)}\mu_{{\boldsymbol{m}}(j)}\bigg)^{1/2}=Ch_{Q}h_{-n}\mathcal{H}^{d}(\Gamma).$
@@
The proposed quadrature formulas require the values of $\mu_{{\boldsymbol{m}}}=\mathcal{H}^{d}(\Gamma_{\boldsymbol{m}})$, which are easily computed in terms of $\mathcal{H}^{d}(\Gamma)$ as $\mu_{{\boldsymbol{m}}}=\rho_{m_{1}}^{d}\cdots\rho_{m_{\ell}}^{d}\mathcal{H}^{d}(\Gamma)$, for all ${\boldsymbol{m}}\in I_{\mathbb{N}}$. For $n=1$, the Hausdorff measure of a class of Cantor sets is shown to be $\mathcal{H}^{d}(\Gamma)=1$ in [23, Thm. 1.14–1.15]; see, e.g., [47] for more recent related results. However, for $n>1$ the exact value of the Hausdorff measure of even the simplest IFS attractors is known only for $d\leq 1$ (see, e.g., [46]), i.e. for the cases that are not relevant for scattering problems (recall, as discussed above Lemma 4.4, that $H^{-1/2}_{\Gamma}=\{0\}$ and $\phi=0$ for $d\leq n-1$). Thus, the entries of the Galerkin matrix (60) and right-hand side vector (61) in our Hausdorff BEM can be computed only up to unknown factors that are powers of the unknown measure $\mathcal{H}^{d}(\Gamma)$, so that the same is true for the vector $\vec{c}$ that is the solution of the Galerkin linear system (57). To be precise, we can’t compute $\vec{c}$, but we can (even if we don’t know $\mathcal{H}^{d}(\Gamma)$) compute ${\vec{c}}^{\,\star}=(c_{1}^{\star},\ldots,c_{N}^{\star})^{T}$, related to $\vec{c}$ by ${\vec{c}}^{\,\star}=(\mathcal{H}^{d}(\Gamma))^{1/2}\vec{c}$.
Surprisingly, this inability to compute $\vec{c}$ does not affect in the slightest the applicability of our Hausdorff BEM. Indeed, it is enough to know $\vec{c}$$^{\,\star}$ (whose value is independent of $\mathcal{H}^{d}(\Gamma)$) to compute the BEM solution $\phi_{N}=\sum_{j=1}^{N}c_{j}e^{j}\in H_{\Gamma}^{-1/2}$. For recall that we know $\phi_{N}$ (as an element of the set of distributions $H_{\Gamma}^{-1/2}$) if we can compute the duality pairing $\langle\varphi,\phi_{N}\rangle_{H^{1/2}(\mathbb{R}^{n})\times H^{-1/2}(\mathbb{R}^{n})}$ for every $\varphi\in H^{1/2}(\mathbb{R}^{n})$. Further, noting that $\operatorname{supp}(f^{j})=\Gamma_{{\boldsymbol{m}}(j)}$ and that $\mu_{\boldsymbol{m}}=\mathcal{H}^{d}(\Gamma_{\boldsymbol{m}})=(\operatorname{diam}(\Gamma_{\boldsymbol{m}})/\operatorname{diam}(\Gamma))^{d}\,\mathcal{H}^{d}(\Gamma)$, we see, using (15), that
$\langle\varphi,\phi_{N}\rangle_{H^{1/2}(\mathbb{R}^{n})\times H^{-1/2}(\mathbb{R}^{n})}=\sum_{j=1}^{N}\overline{c_{j}}\int_{\Gamma}f^{j}\mathrm{tr}_{\Gamma}\varphi\,\mathrm{d}\mathcal{H}^{d}=\sum_{j=1}^{N}\overline{c^{\star}_{j}}\left(\textstyle{\frac{\operatorname{diam}(\Gamma)}{\operatorname{diam}(\Gamma_{{\boldsymbol{m}}(j)})}}\right)^{d/2}\frac{1}{\mathcal{H}^{d}(\Gamma)}\int_{\Gamma_{{\boldsymbol{m}}(j)}}\mathrm{tr}_{\Gamma}\varphi\,\mathrm{d}\mathcal{H}^{d},$
for every $\varphi\in H^{1/2}(\mathbb{R}^{n})$, and the right hand side of the above is independent of the value of $\mathcal{H}^{d}(\Gamma)$.
A simple implementation technique which avoids working with an unknown value for $\mathcal{H}^{d}(\Gamma)$ and produces the correct solution $\phi_{N}$ is just to set $\mathcal{H}^{d}(\Gamma)=1$ in all calculations. This is equivalent to introducing a “normalised Hausdorff measure” $\mathcal{H}^{d}_{\star}(\cdot):=\mathcal{H}^{d}(\cdot)/\mathcal{H}^{d}(\Gamma)$, so that $\mathcal{H}^{d}_{\star}(\Gamma)=1$, and using it throughout in the BEM in place of $\mathcal{H}^{d}(\cdot)$.
@@
Now $\phi^{Q}_{N}\in V_{N}$ is the solution of $a^{Q}(\phi^{Q}_{N},\psi_{N})=B^{Q}(\psi_{N})$, $\forall\psi_{N}\in V_{N}$. Strang’s first lemma (e.g. [5, III.1.1]), applied in the special case that $V=S_{h}$ (in the notation of [5, III.1.1]), combined with (109) and (110), gives
@@
The quadrature rule defined by (94) uses the same maximum mesh width $h_{Q}$ for all off-diagonal elements $A_{ij}$, $i\neq j$. Since the magnitude of the integrand $\Phi(x,y)$ and its derivatives blows up as one approaches the diagonal $x=y$, and decays away from it, it is possible to save computational effort, while maintaining the error bounds (104) and (105), and hence the error estimates (112) and (118). To achieve this we increase the quadrature mesh width (and hence potentially decrease the number of quadrature points) for the computation of $A_{ij}$ when the elements $\Gamma_{{\boldsymbol{m}}(i)}$ and $\Gamma_{{\boldsymbol{m}}(j)}$ are sufficiently well-separated spatially. In more detail, from [25, Prop 5.2] we have that, for $i\neq j$, the error in the quadrature approximation (94), with $h_{Q}$ replaced by a local quadrature mesh width $h_{Q,i,j}$, satisfies

### v2

@@ 4.1 The BIE on $d$-sets in trace spaces
$\langle\mathbb{S}\Psi,\widetilde{\Psi}\rangle_{\mathbb{H}^{t_{d}}(\Gamma)\times\mathbb{H}^{-t_{d}}(\Gamma)}=-\langle\mathrm{tr}_{\Gamma}g,\widetilde{\Psi}\rangle_{\mathbb{H}^{t_{d}}(\Gamma)\times\mathbb{H}^{-t_{d}}(\Gamma)},\qquad\mbox{ for all }\widetilde{\Psi}\in\mathbb{H}^{-t_{d}}(\Gamma),$ (47)
@@
The definition and mapping properties of $\mathrm{tr}_{\Gamma}$ and $\mathrm{tr}_{\Gamma}^{*}$, noted in §2.4, combined with the representation (48), enable us to extend the domain of $\mathbb{S}$ to $\mathbb{H}^{-t}(\Gamma)$, for $t_{d}<t<2t_{d}$, or restrict it to $\mathbb{H}^{-t}(\Gamma)$, for $0<t<t_{d}$, as stated in the following key result.
@@
Let $\Omega\subset\Gamma_{\infty}$ be any bounded open set containing $\Gamma$, so that (see §2.4) $H^{s}_{\Gamma}$ is a closed subspace of $\widetilde{H}^{s}(\Omega)$ for every $s\in\mathbb{R}$. The claimed mapping property of $\mathbb{S}$ follows from (48) since $\mathrm{tr}_{\Gamma}:H^{s}(\mathbb{R}^{n})\to\mathbb{H}^{t}(\Gamma)$ and $\mathrm{tr}_{\Gamma}^{*}:\mathbb{H}^{-t}(\Gamma)\to H^{-s}_{\Gamma}\subset\widetilde{H}^{-s}(\Omega)$ are continuous for s>(nd)/2s>(n-d)/2 (i.e. $t>0$), with $s$ and $t$ related by (9), and since the mapping $\phi\mapsto\gamma^{\pm}((\sigma\mathcal{S}\phi)|_{U^{\pm}}):\widetilde{H}^{s}(\Omega)\to H^{s+1}(\Gamma_{\infty})$ is continuous for $s\in\mathbb{R}$ (e.g., [14, Thm 1.6]). ∎
@@
We recall the following result from [39, Prop. 4.7], which quotes [45]. Suppose $E_{j}$ and $F_{j}$ are Banach spaces, $E_{1}\subset E_{0}$ and $F_{1}\subset F_{0}$ with continuous embeddings, and $T:E_{j}\to F_{j}$ is a bounded linear operator for $j=0,1$. Let $E_{\theta}=(E_{0},E_{1})_{\theta}$ be defined by complex interpolation, and similarly $F_{\theta}$, so $T:E_{\theta}\to F_{\theta}$ and is bounded, for $0<\theta<1$. Assume that $T:E_{\theta_{0}}\to F_{\theta_{0}}$ is invertible for some $\theta_{0}\in(0,1)$. Then $T:E_{\theta}\to F_{\theta}$ is invertible for $\theta$ in a neighbourhood of $\theta_{0}$.
@@ 5 The Hausdorff BEM
$a(\phi_{N},\psi_{N})=-\langle g,\psi_{N}\rangle_{H^{1/2}(\Gamma_{\infty})\times H^{-1/2}(\Gamma_{\infty})},\qquad\mbox{ for all }\psi_{N}\in V_{N}.$ (54)
@@
The well-posedness of (54) follows from the Lax–Milgram lemma and the continuity and coercivity of $)S:H^{-1/2}_{\Gamma}\to(\widetilde{H}{}^{1/2}(\Gamma^{c}$$)^{\perp}$. Furthermore, by Céa’s lemma (e.g., [47, Theorem 8.1]) we have the following quasi-optimality estimate, where $C_{a},\alpha>0$ are as in (39):
@@ 5.3 Inverse estimates and conditioning
The first claim of the following corollary follows immediately from the above result. The second (cf. [25, Thm. 1.4], [22], [20, Thm. 3.2.4]) combines our earlier results with standard Aubin-Nitsche lemma arguments.
@@ 5.4 Numerical quadrature and fully discrete error estimates
For $i\neq j$ the integrand in (91) is continuous, because, by assumption, the components of $\Gamma$ are disjoint. In that case, to evaluate (91) we use the composite barycentre rule of [30, Defn 3.5]. In our context, this means partitioning the BEM elements $\Gamma_{{\boldsymbol{m}}(i)}$ and $\Gamma_{{\boldsymbol{m}}(j)}$, which have diameter approximately equal to $h$, into a union of self-similar subsets of a possibly smaller approximate diameter $0<h_{Q}\leq h$, writing the integral in (91) as a sum of integrals over the Cartesian products of these subsets, and approximating these integrals by a one-point barycentre rule. Specifically, we approximate
@@
is the barycentre of $\Gamma_{\boldsymbol{n}}$ with respect to the measure $\mathcal{H}^{d}$. (Note that $x_{\boldsymbol{n}}$ is not necessarily an element of $\Gamma_{\boldsymbol{n}}$.) The similarities can be written as $s_{m}(x)=\rho_{m}A_{m}x+v_{m}$, for some $v_{m}\in\mathbb{R}^{n}$ and some orthogonal $A_{m}\in\mathbb{R}^{n\times n}$ ($A_{m}=\pm 1$ in the case $n=1$), and then, writing $\boldsymbol{n}=(n_{1},\ldots,n_{\ell})$, we find (see [30, Prop. 3.3]) that
@@
For $i=j$ the integral (91) is singular. To evaluate it we adopt the singularity subtraction approach of [30, §5], writing
@@
The integrals appearing in (95) all have continous integrands and can be evaluated using the composite barycentre rule (as in [30, Eqn (47)]), viz.
@@
Assuming the data $g$ is sufficiently smooth, the right-hand side entries (59) are regular single integrals and can be evaluated by a single integral version of the composite barycentre rules described for double integrals above, as in [30, Defn 3.1]). Explicitly, we use the approximation
@@
The following quadrature error estimates follow from results in [30], specifically [30, Prop. 5.2, Thms 3.6(iii), 5.7 & 5.11]. Following the terminology of [30], we say an IFS attractor $\Gamma$ is hull-disjoint if the convex hulls $\mathrm{Hull}(\Gamma_{1}),\ldots,\mathrm{Hull}(\Gamma_{M})$ are disjoint, which holds if and only if the OSC holds for some open set $O\supset\mathrm{Hull}(\Gamma)$ (cf. Lemma 2.5, the proof of which works in the same way for hull-disjointness as for disjointness). The assumption of hull-disjointness permits analysis of the singular quadrature rules in [30] by Taylor expansion. (The point here is that while the barycentre $x_{{\boldsymbol{m}}}$ of a component $\Gamma_{{\boldsymbol{m}}}$ may not lie in $\Gamma_{{\boldsymbol{m}}}$, it must lie in $\mathrm{Hull}(\Gamma_{{\boldsymbol{m}}}$).) However, numerical experiments suggest that hull-disjointness is not essential for the applicability of the quadrature rules in [30] (see [30, §6]). We also suspect that the estimates (103) and (102) may not be sharp in their $h$-dependence — see [30, Rem. 5.10] and the discussion around [30, Fig. 7]. In Remark 5.19 we describe modifications to our quadrature rule which may improve its efficiency.
@@
the first estimate in (100) follows from [30, Thm 3.6(iii)], and the second then follows from the fact that $\sum_{i=1}^{N}\mu_{{\boldsymbol{m}}(i)}=\mathcal{H}^{d}(\Gamma)$. For (ii), given $\psi_{N}\in V_{N}$ let ${\vec{\psi}}=(({\vec{\psi}})_{1},\ldots,({\vec{\psi}})_{N})^{T}\in\mathbb{C}^{N}$ denote the coefficient vector of its expansion with respect to the basis $(e^{1},\ldots,e^{N})$. By the orthonormality of the basis $(f^{1},\ldots,f^{N})$ in $\mathbb{L}_{2}(\Gamma)$, the inverse estimate (84) and the isometry property (14), given $0<t<1$ we can bound
@@
Then, to prove (101), using (63), (99), [30, Thm 3.6(iii)], (104), and the Cauchy-Schwarz inequality,
@@
For (iii), the first estimates in (102) and (103) follow from [30, Prop 5.2] for the off-diagonal terms and [30, Thms 5.7 & 5.11] for the diagonal terms. The factors of $h^{-(n+1)}$ and $h^{-n}$ come from the fact that, in the notation of [30, Thms 5.7 & 5.11], $R_{\Gamma,\mathrm{Hull}}$ equals $\operatorname{diam}(\Gamma)$ times a constant independent of $\operatorname{diam}(\Gamma)$. We note that, since we are allowing $C$ to be $k$-dependent, we can use the $k\operatorname{diam}(\Gamma)\leq c_{\rm osc}$” estimates in [30, Thms 5.7 & 5.11], since, given $k>0$, the constant $c_{\rm osc}$ can be chosen as large as is required. To derive the second estimate in (102) (a similar argument gives the second estimate in (103)), note that, with $\|\cdot\|_{\mathrm{F}}$ denoting the Frobenius norm,
$\|A-A^{Q}\|_{2}\leq\|A-A^{Q}\|_{\mathrm{F}}\leq Ch_{Q}h^{-n}\bigg(\sum_{i=1}^{N}\sum_{j=1}^{N}\mu_{{\boldsymbol{m}}(i)}\mu_{{\boldsymbol{m}}(j)}\bigg)^{1/2}=Ch_{Q}h^{-n}\mathcal{H}^{d}(\Gamma).$
@@
The proposed quadrature formulas require the values of $\mu_{{\boldsymbol{m}}}=\mathcal{H}^{d}(\Gamma_{\boldsymbol{m}})$, which are easily computed in terms of $\mathcal{H}^{d}(\Gamma)$ as $\mu_{{\boldsymbol{m}}}=\rho_{m_{1}}^{d}\cdots\rho_{m_{\ell}}^{d}\mathcal{H}^{d}(\Gamma)$, for all ${\boldsymbol{m}}\in I_{\mathbb{N}}$. For $n=1$, the Hausdorff measure of a class of Cantor sets is shown to be $\mathcal{H}^{d}(\Gamma)=1$ in [27, Thm. 1.14–1.15]; see, e.g., [54] for more recent related results. However, for $n>1$ the exact value of the Hausdorff measure of even the simplest IFS attractors is known only for $d\leq 1$ (see, e.g., [53]), i.e. for the cases that are not relevant for scattering problems (recall, as discussed above Lemma 4.4, that $H^{-1/2}_{\Gamma}=\{0\}$ and $\phi=0$ for $d\leq n-1$). A simple implementation technique which avoids working with an unknown value for $\mathcal{H}^{d}(\Gamma)$ and produces the correct solution $\phi_{N}$ is just to set $\mathcal{H}^{d}(\Gamma)=1$ in all calculations. This is equivalent to introducing a “normalised Hausdorff measure” $\mathcal{H}^{d}_{\star}(\cdot):=\mathcal{H}^{d}(\cdot)/\mathcal{H}^{d}(\Gamma)$, so that $\mathcal{H}^{d}_{\star}(\Gamma)=1$, and using it throughout in the BEM in place of $\mathcal{H}^{d}(\cdot)$.
@@
Now $\phi^{Q}_{N}\in V_{N}$ is the solution of $a^{Q}(\phi^{Q}_{N},\psi_{N})=B^{Q}(\psi_{N})$, $\forall\psi_{N}\in V_{N}$. Strang’s first lemma (e.g. [6, III.1.1]), applied in the special case that $V=S_{h}$ (in the notation of [6, III.1.1]), combined with (107) and (108), gives
@@
The quadrature rule defined by (92) uses the same maximum mesh width $h_{Q}$ for all off-diagonal elements $A_{ij}$, $i\neq j$. Since the magnitude of the integrand $\Phi(x,y)$ and its derivatives blows up as one approaches the diagonal $x=y$, and decays away from it, it is possible to save computational effort, while maintaining the error bounds (102) and (103), and hence the error estimates (110) and (116). To achieve this we increase the quadrature mesh width (and hence potentially decrease the number of quadrature points) for the computation of $A_{ij}$ when the elements $\Gamma_{{\boldsymbol{m}}(i)}$ and $\Gamma_{{\boldsymbol{m}}(j)}$ are sufficiently well-separated spatially. In more detail, from [30, Prop 5.2] we have that, for $i\neq j$, the error in the quadrature approximation (92), with $h_{Q}$ replaced by a local quadrature mesh width $h_{Q,i,j}$, satisfies